in an evening, a sporting goods store sold twice as monay T-shirts as shorts. t-shirts are $9 each, and shorts are $14 each. The total amount of money taken in for both items was $256. Find the number of each that was sold ( using and method ).
Answer by ptaylor(1894)


Let x=number of shorts sold
(1) Amount of money taken in from shorts=14x
Then 2x=number of T-shirts sold
(2) Amount of money taken in from T-shirts=9*2x or 18x
Now we are told that (1)+(2)=$256. So our equation to solve is:
14x+18x=256 collect like terms
32x=256 divide both sides by 32
x=8----------------------------number of shorts sold
2x=2*8=16-----------------------number of T-shirts sold
CK
8*$14+16*9=$256
$112+$144=$256
$256=$256

Hope this helps---ptaylor